### v2

@@ Abstract
Coherent elastic neutrino-nucleus scattering (CENS) provides a unique probe for neutrino properties Beyond the Standard Model (BSM) physics. REactor neutrino LIquid xenon Coherent Scattering experiment (RELICS), a proposed reactor neutrino program using liquid xenon time projection chamber (LXeTPC) technology, aims to investigate the CENS process of antineutrinos off xenon atomic nuclei. In this work, the design of the experiment is studied and optimized based on Monte Carlo (MC) simulations. To achieve a sufficiently low energy threshold for CENS detection, an ionization-only analysis channel is adopted for RELICS. A high emission rate of delayed electrons after a big ionization signal is the major background, leading to an analysis threshold of 120 photo-electrons in the CENS search. The second largest background, nuclear recoils induced by cosmic-ray neutrons, is suppressed via a passive water shield. The physics potential of RELICS is explored with a 32 exposure at a baseline of 25 m from a reactor core with a 3 GW thermal power. In an energy range of 120 to 240 PE, we expect 4902.4 CENS and 1318.4 background events. The sensitivity of RELICS to the weak mixing angle is investigated at a low momentum transfer. Our study shows that RELICS can further improve the constraints on the non-standard neutrino interaction (NSI) compared to the current best results.
@@
Coherent elastic neutrino-nucleus scattering (CENS) is a Standard Model (SM) weak neutral current process proposed in 1974 [1]. The neutrino treats the nucleus as a single particle during scattering with a cross-section proportional to the nucleon number squared. Despite this process having a relatively large cross-section, the detection is challenging due to the ultra-low deposited energy. This process had evaded experimental discoveries for decades until 2017 when it was first observed by the COHERENT collaboration with neutrinos produced by the Spallation Neutron Source (SNS) at Oak Ridge National Laboratory scattering off a CsI scintillator [2], and later confirmed with a liquid argon detector [3] in 2020. These measurements of CENS offer new probes for diverse physical phenomena.
@@
CENS can provide new constraints on the Standard Model physics and the physics Beyond it (BSM). The weak mixing angle is related to the CENS cross-section [4]. Any deviation from the SM prediction in measurement can be an indication of new physics. CENS is also a valuable tool to constrain nonstandard neutrino interactions (NSI) [5], to search for sterile neutrinos [6], to verify the existence of non-trivial neutrino electromagnetic properties [7], to study the nuclear structure [8] and astrophysics [9]. In addition, the CENS signal from the solar or atmospheric neutrinos is expected to be an irreducible background for the next-generation dark matter direct detection experiments, such as PandaX-xT [10] and XLZD [11]. A precise measurement of this process, especially one with a xenon target, can provide valuable information for dark matter experiments digging into the so-called “neutrino floor/fog” [12, 13].
@@
The observation of CENS by COHERENT collaboration inspired the interest in CENS experimental study. While the neutrino beam has been the primary source for the CENS discovery, it is limited by the total flux delivered and the availability of the facility. Therefore, nuclear reactors, with large anti-neutrino fluxes delivered with typical energies of a few MeV, which is the so-called “fully coherent regime” of CENS , have been proposed as attractive alternative facilities for the CENS study. Due to the considerably low recoil energy and complex background, the reactor CENS is harder to detect. Many efforts are underway with reactor neutrinos, such as [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Detecting reactor CENS will provide a wealth of physics opportunities and potential applications such as nuclear safeguards [24].
@@
RELICS is a reactor neutrino experiment using the LXeTPC technology. LXeTPC is one of the leading technologies for rare event searches, such as detecting dark matter and neutrinoless double beta decay [11, 25, 26, 27, 28, 29]. Benefiting from its ultra-low threshold by using ionization-only channel [30, 31] and reduced background, the LXeTPC is an excellent approach for reactor CENS detection. The RELICS experiment is proposed to be located near the Sanmen Nuclear Power Plant in Zhejiang province, China. At a 25 m baseline to the 3 GW reactor core, the average neutrino flux available for RELICS is 1013cm-2s-1.
@@ II.1 RELICS detector
The LXeTPC technology measures the energy deposited from particle interactions by simultaneously detecting the prompt scintillation light (S1) and delayed ionization charge (S2) [32]. The S1 is produced by the direct excitation of xenon atoms and the recombination of electron-ion pairs in LXe. The S2 is proportional to the number of liberated electrons extracted into gas xenon (GXe), which are then amplified via electroluminescence. Both S1 and S2 are recorded by photosensors, usually photomultiplier tubes (PMTs), on top and bottom arrays of the TPC. This allows a three-dimensional (3D) vertex reconstruction with a sub-centimeter precision, and the S2/S1 ratio can be used to discriminate between nuclear recoil (NR) and electronic recoil (ER) for background reduction. Typically, the energy threshold of LXeTPC is limited by the S1 signal in 23 photo-electrons (PEs). The inherent S2 amplification provides an alternative approach to decrease the energy threshold by utilizing the ionization-only channel, usually called the “S2-only” analysis. This approach can reduce the energy threshold to a few ionization electrons (equivalent NR energy below 1 keVnr) by sacrificing the NR/ER discrimination and depth information, which, in the case of CENS detection, can be largely compensated by the high signal-to-background ratio profited from the intense reactor neutrino flux.
@@ III.1 Liquid xenon response
Particle interaction in LXe will produce photons and electrons. The NEST model can characterize this process [36] after considering the fluctuations in scintillation, ionization, electron-ion recombination, and their drift field dependence. In RelicsAPT, the Ly and Qy from NEST (v2.3.6) at 500 V/cm drift field are used with no uncertainty assigned for NRs larger than 3 keVnr. To quantify the uncertainty of light and charge yields for CENS events, the Ly and Qy bands in XENON1T solar neutrino search [37] are adopted for NRs less than 3 keVnr. The bands of Ly and Qy are scaled to make the center curve connected to NEST at 3 keVnr. The final Ly and Qy of NRs applied in RelicsAPT are shown in figure 3. For ERs, Ly and Qy from NEST (v2.3.6) are used without uncertainties. In the 4 LXe veto region, both Ly and Qy are modeled with zero electric fields.
@@ III.2 Detector related effects
To explore the capability of the RELICS detector on the ionization-only detection channel, the light collection efficiency (LCE) of S2 is simulated by RelicsSim with the optical parameters listed in Tab. 1. The LCE map is shown in figure 4. An averaged value of 30 PE/e- is estimated based on the S2 LCE map, the properties of the PMTs, such as , , and that also listed in Tab. 1, and a 100% electron extraction efficiency on the liquid surface. An electron lifetime of 1.2 s is applied for S2 estimation considering the probability of electrons loss to electronegative impurities during their drift.
@@ IV Expected CENS events
where is Fermi’s coupling constant, is the weak nuclear charge, is the ground state elastic form factor, is the mass of the target nucleus, is the nuclear recoil energy while the incident neutrino energy. is required to ensure coherence, where is the nuclear radius. Normally, this can be satisfied when MeV.
@@
For the reactor neutrinos with energy above 2 MeV, the Huber-Mueller model[45, 46] is adopted for the spectra calculation while P. Vogel’s theory [47] is used for below 2 MeV region, which is rarely searched due to the 1.8 MeV energy threshold of inverse beta decay reaction. We assume a fuel mixture of 235U (56.1%), 238U (7.6%), 239Pu (30.7%), and 241Pu (5.6%), the total neutrino flux can be estimated with a 3 GW reactor core at 25m distance, and then the CENS event rate can be obtained accordingly as shown in figure 5.
@@ V Background contributions
The complexity of the backgrounds in reactor CENS detection is one of the major reasons it escaped the searches for several decades. To overcome this challenge, dedicated background simulations have been conducted to optimize the design of the shielding system for RELICS, using RelicsSim and RelicsAPT. The background contributions can be divided into five categories according to their origin, such as (V.1) cosmic-ray neutrons, (V.2) cosmic muons, (V.3) detector materials, and (V.4) internal background from LXe. The last one is (V.5) delayed electrons, which have been observed by large LXeTPC searching for dark matter and should be treated carefully.
This section describes all the background components and is summarized in (V.6). The following criteria are applied to MC data to select the background events that can mimic CENS signals:
@@ V.1 Background induced by cosmic-ray neutrons
The cosmic rays that travel through the atmosphere can produce a variety of radiation particles, such as neutrons, protons, pions, etc., through spallation reaction on nitrogen, oxygen, etc [48, 49]. Among them, the high-energy cosmic-ray neutrons (CRNs) can easily pass through the shield and deposit energy in the RELICS detector. The main purpose of the water shield design is to block the CRNs. The size of the water tank is optimized based on the MC simulation. The CRN flux is calculated using the Cosmic-Ray Shower Generator (CRY) source [50]. The energy spectrum of the CRNs from CRY is shown in figure 6. Analyze the MC data shows that only CRNs with energy greater than 103 MeV can deposit energy in LXe, which indicates that the neutrons from the nuclear reactor and laboratory environment normally with energy less than 10 MeV [51] that also shows in figure 6, can not contribute to the backgrounds.
After processing the MC data, and passing the LXe-veto, single-scatter and FV selections, the CRNs-induced NR background rate is estimated to be (7.7 0.7) in keVnr region-of-interest (ROI) of CENS searching, with the recoil energy spectrum shows in figure 7 (top). In addition, the secondaries of CRNs will also induce ER background in LXe, with a rate of (3.9 0.3), which is shown in figure 7 (bottom).
@@ V.2 Background induced by cosmic muons
The RELICS detector will operate without overburden. Hence, special attention has to be paid to the muon-related background. Even though muons can be tagged with 99% efficiency by the PS veto detector, the neutron background produced by muon interacting with the detector materials is still one of the main background components. Dedicated simulations are performed to evaluate the muon-induced backgrounds. The energy and angular distributions of muons follow the Shukla model [52]. The energy deposition in LXe that is caused by muon-induced fast neutron and related gammas are recorded by RelicsSim, and then filtered with LXe-veto, single-scatter and FV selections. Assuming a 99% muon tagging efficiency in 300 veto time, the muon-induced NRs and ERs are (0.08 0.01) in keVnr and (28.0 2) in keVer, respectively. The corresponding spectra are shown in figure 7.
@@ V.3 Neutrons and gammas from the detector materials
S2 width selection: As shown in figure 8, a large amount of background events are located close to the liquid-gas interface. These backgrounds are mainly from the radioactivities of PMTs and can not be suppressed without the selection in “z” in an “S2-only” analysis. As demonstrated in [30], the pulse shape of S2 signals can discriminate background happening at the liquid-gas interface to signals distributed uniformly in depth “z”. A toy MC was conducted to consider the drift and diffusion properties of the electron clouds in LXe to optimize such a selection. We simulate S2s and their pulse shape parameters with “z” distributions shown in figure 8 with energy in the CENS ROI. The S2s are simulated via RelicsAPT where the NEST model is incorporated. The S2 pulse shape is simulated using the electron diffusion model , where is the longitudinal diffusion coefficient, is the electron drift velocity, is the width of single electron waveform [54]. In our simulation. = 16 cm2/s, = 0.174 cm/s, and = 0.19s.
@@ V.4.2 Cosmogenic isotopes and
The RELICS experiment, which is to be operated without overburden, is unavoidable under cosmic radiation. We consider cosmogenic radioactive isotopes through the activation and spallation of xenon atoms. In particular, and are known cosmogenic backgrounds in LXeTPCs [57, 27] due to their characteristic x-rays or Auger electrons through electron capture process and their relatively long half-life of 36.4 and 35.0 days, respectively.
Natural xenon does not contain . can be produced from when cosmogenic muons kick off a neutron in the nucleus with a cross-section of (2.74 0.4) cm2. Additionally, we consider another production channel through the neutron capture on with an upper limit cross-section of 0.7cm2 at 95% C.L [58]. The production rate of in LXe is calculated based on its production cross-section, the muon’s fluxes and energy spectrum, and thermal neutrons. The decays to via electron capture (EC) process, with a half-life of 36.4 days. We assume that the production and decay of reach equilibrium before deploying the RELICS detector inside the shielding. The estimated equilibrium decay rate is 9.29 atoms/day/kg. The shielding system of RELICS can attenuate neutrons efficiently, and the equilibrium decay rate can decrease to 0.79 atoms/day/kg after 300 days.
@@
We consider three possible mechanisms for production in xenon, following the approaches in [59]. Firstly, the extremely low content of in the atmosphere, which is generated by spallation of through (n, 4n), neutron capture of , or cosmic bombardment of calcium-containing soils, via (n, ). This mechanism only needs to be considered in the context of air leaks of an LXe detector, which is strictly limited to below 0.1 liter per year to ensure the purity level of LXe. Secondly, the spallation of detector materials, especially iron in steel. However, its contribution is greatly suppressed due to the slow diffusion rate of argon in steel. Lastly, the direct spallation of xenon atoms by high-energy cosmogenic protons and neutrons. Only the last possibility is considered in this work. The semi-empirical formula by Silberberg and Tsao was used to calculate the spallation cross-section. The CRY is used to calculate the proton spectrum. Gordon’s neutron spectrum [60] is adopted for cosmogenic neutron estimation. The production rate of this process is estimated to be 0.024 atoms/kg/day with an uncertainty of 100%. The produced decays to the ground state of via EC and subsequently deposits energy via x-rays or Auger electrons. The total energy (relative intensity) for K, L, and M shells are 2.82 keV (90.2%), 0.270 keV (8.9%), and 0.018 keV (0.9%), respectively. As a results, the background rate of in the CENS ROI is below kg-1day-1 and completely negligible for CENS search.

### v3

@@ Abstract
Coherent elastic neutrino-nucleus scattering (CENS) provides a unique probe for neutrino properties Beyond the Standard Model (BSM) physics. REactor neutrino LIquid xenon Coherent Scattering experiment (RELICS), a proposed reactor neutrino program using liquid xenon time projection chamber (LXeTPC) technology, aims to investigate the CENS process of antineutrinos off xenon atomic nuclei. In this work, the design of the experiment is studied and optimized based on Monte Carlo (MC) simulations. To achieve a sufficiently low energy threshold for CENS detection, an ionization-only analysis channel is adopted for RELICS. A high emission rate of delayed electrons after a big ionization signal is the major background, leading to an analysis threshold of 120 photo-electrons in the CENS search. The second largest background, nuclear recoils induced by cosmic-ray neutrons, is suppressed via a passive water shield. The physics potential of RELICS is explored with a 32 exposure at a baseline of 25 m from a reactor core with a 3 GW thermal power. In an energy range of 120 to 300 PE, corresponding to an average nuclear recoil from 0.63 to 1.36 keV considering the liquid xenon response and detector-related effect, we expect 4639.7 CENS and 1687.8 background events. The sensitivity of RELICS to the weak mixing angle is investigated at a low momentum transfer. Our study shows that RELICS can further improve the constraints on the non-standard neutrino interaction (NSI) compared to the current best results.
@@
Coherent elastic neutrino-nucleus scattering (CENS) is a Standard Model (SM) weak neutral current process postulated in 1974 [1]. The incoming neutrino sees the nucleus as a whole during scattering with a cross-section proportional to the nucleon number squared in the full coherency regime when the neutrino energy is smaller than 10 MeV. Despite this process having a relatively large cross-section, the detection is challenging due to the ultra-low deposited energy in keV or sub-keV nuclear recoil scale. This process had evaded experimental discoveries for decades until 2017 when it was first observed by the COHERENT collaboration with neutrinos produced by the Spallation Neutron Source (SNS) at Oak Ridge National Laboratory scattering off a CsI scintillator [2], and later confirmed with a liquid argon detector [3] in 2020. These measurements of CENS offer new probes for diverse physical phenomena.
@@
CENS can provide new constraints on the SM physics and the physics Beyond it (BSM). The weak mixing angle is related to the CENS cross-section [4]. Any deviation from the SM prediction in measurement can be an indication of new physics. CENS is also a valuable tool to constrain nonstandard neutrino interactions (NSI) [5], to search for sterile neutrinos [6], to verify the existence of non-trivial neutrino electromagnetic properties [7], to study the nuclear structure [8] and understand the core-collapse supernova explosions [9]. In addition, the CENS signal from the solar or atmospheric neutrinos is expected to be an irreducible background for the next-generation dark matter direct detection experiments, such as PandaX-xT [10] and XLZD [11] using xenon as detection target [12, 13]. A precise measurement of this process, especially one with a xenon target, can provide valuable information for dark matter experiments digging into the so-called “neutrino floor/fog” [14, 15, 16].
@@
The observation of CENS by the COHERENT collaboration opened the way to CENS experimental studies. While the neutrino beam has been the primary source for the CENS discovery, the anti-neutrino from nuclear reactors provides an alternative chance for CENS detection. Therefore, the nuclear reactors with large anti-neutrino fluxes delivered with typical energies of a few MeV, which is in the full coherency regime of CENS, have been proposed as attractive complementary facilities for several CENS experiments. Due to the considerably low recoil energy and complex background, the reactor CENS is harder to detect. Many efforts are underway with reactor neutrinos, such as [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Benefiting from the higher flux and lower energy of reactor neutrinos compared with the SNS neutrino beam, the detection of reactor CENS can provide a complementary sensitivity in most fields of the physical phenomena mentioned above through new couplings [29, 30], and also potential applications such as nuclear safeguards [31].
@@
In this context, RELICS is a reactor neutrino experiment using liquid xenon time projection chamber (LXeTPC) technology. LXeTPC is one of the leading technologies for rare event searches, such as detecting dark matter and neutrinoless double beta decay [11, 32, 33, 34, 35, 36]. Benefiting from its ultra-low threshold by using ionization-only channel [37, 38] and reduced background, the LXeTPC is an excellent approach for reactor CENS detection. The RELICS experiment is proposed to be located near the Sanmen Nuclear Power Plant in Zhejiang province, China. At a 25 m baseline to the 3 GW reactor core, the average neutrino flux available for RELICS is 1013cm-2s-1.
@@ II.1 RELICS detector
The LXeTPC technology consists in measuring the energy deposited from particle interactions by simultaneously detecting the prompt scintillation light (S1) and delayed ionization charge (S2) [39]. The S1 is produced by the direct excitation of xenon atoms and the recombination of electron-ion pairs in LXe. The S2 is proportional to the number of liberated electrons extracted into gas xenon (GXe), which are then amplified via electroluminescence. Both S1 and S2 are recorded by photosensors, usually photomultiplier tubes (PMTs), on top and bottom arrays of the TPC. This allows a three-dimensional (3D) vertex reconstruction with a sub-centimeter precision, and the S2/S1 ratio can be used to discriminate between nuclear recoil (NR) and electronic recoil (ER) for background reduction. Typically, the energy threshold of LXeTPC is limited by the S1 signal in the range of (2-3) photo-electrons (PEs). The inherent S2 amplification provides an alternative approach to decrease the energy threshold by utilizing the ionization-only channel, usually called the “S2-only” analysis. This approach can reduce the energy threshold to a few ionization electrons (equivalent NR energy below 1 keVnr) by sacrificing the NR/ER discrimination and depth information, which, in the case of CENS detection, can be largely compensated by the high signal-to-background ratio profited from the intense reactor neutrino flux.
@@ III.1 Liquid xenon response
Particle interaction in LXe will produce photons and electrons. The NEST model can characterize this process [43] after considering the fluctuations in scintillation, ionization, electron-ion recombination, and their drift field dependence. In RelicsAPT, the Ly and Qy from NEST (v2.3.6) at 500 V/cm drift field are used with no uncertainty assigned for NRs larger than 3 keVnr. To quantify the uncertainty of light and charge yields for CENS events, the Ly and Qy bands in XENON1T solar neutrino search [44] are adopted for NRs less than 3 keVnr. Ly and Qy bands are scaled to make the center curve connected to NEST at 3 keVnr. The final Ly and Qy of NRs applied in RelicsAPT are shown in figure 3. These light and charge yield responses and their uncertainties comparable to the XENON1T model are chosen as a benchmark model in this analysis to illustrate its impact on the sensitivity to CENS and new physics beyond the Standard Model. For ERs, Ly and Qy from NEST (v2.3.6) are used without uncertainties. In the 4 LXe veto region, both Ly and Qy are modeled with zero electric fields.
@@ III.2 Detector related effects
To explore the capability of the RELICS detector on the ionization-only detection channel, the light collection efficiency (LCE) of S2 is simulated by RelicsSim with the optical parameters listed in Tab. 1. The LCE map is shown in figure 4. An averaged value of 30 PE/e- is estimated based on the S2 LCE map, the properties of the PMTs, such as , , and that also listed in Tab. 1, and a 100% electron extraction efficiency on the liquid surface. An electron lifetime of 1.2 s is applied for S2 estimation considering the probability of electrons loss to electronegative impurities during their drift. Under these hypotheses, we assume a 100% trigger efficiency at 4 e- (120 PE) targeted detection threshold for RELICS sensitivity study.
@@ IV Expected CENS events
where is Fermi’s coupling constant, is the weak nuclear charge, is the ground state elastic form factor, is the mass of the target nucleus, is the nuclear recoil energy and the incident neutrino energy. is required to ensure coherence, where is the nuclear radius. This condition can be largely satisfied when MeV.
@@
For the reactor neutrinos with energy above 2 MeV, the Huber-Mueller model[52, 53] is adopted for the spectra calculation while P. Vogel’s calculation [54] is used for below 2 MeV region, which has never been experimentally explored due to the 1.8 MeV energy threshold of inverse beta decay reaction. We assume a fuel mixture of 235U (56.1%), 238U (7.6%), 239Pu (30.7%), and 241Pu (5.6%) [55], the total neutrino flux can be estimated with a 3 GW reactor core at 25m distance, and then the CENS event rate in xenon medium can be obtained accordingly as shown in figure 5 (red). The event rate in RELICS region-of-interest (ROI) of [120, 300] PE is overlaid with (purple) and without (black) signal acceptance, respectively, which will be detailed in Sec. V.
@@ V Background contributions
The complexity of the backgrounds is one of the main challenges of reactor CENS detection. To overcome this challenge, dedicated background simulations have been conducted to optimize the design of the shielding system for RELICS, using RelicsSim and RelicsAPT. The background contributions can be divided into five categories according to their origin, such as (V.1) cosmic-ray neutrons, (V.2) cosmic muons, (V.3) detector materials, and (V.4) internal background from LXe. The last one is (V.5) delayed electrons, which have been observed by large LXeTPC searching for dark matter and should be treated carefully.
This section describes all the background components and is summarized in (V.6). The uncertainties are due to the limited number of simulated exposures. The following criteria are applied to MC data to select the background events that can mimic CENS signals:
@@ V.1 Background induced by cosmic-ray neutrons
The cosmic rays that travel through the atmosphere can produce a variety of radiation particles, such as neutrons, protons, pions, etc., through spallation reaction on nitrogen, oxygen, etc [56, 57]. Among them, the high-energy cosmic-ray neutrons (CRNs) can easily pass through the shield and deposit energy in the RELICS detector. The main purpose of the water shield design is to block the CRNs. The size of the water tank is optimized based on the MC simulation. The CRN flux is calculated using the Cosmic-Ray Shower Generator (CRY) source [58]. Figure 6 shows the energy spectrum of the CRNs from CRY as the blue curve outside the water shield. Analyzing the MC data shows that only CRNs with energy greater than 103 MeV can deposit energy in LXe, which indicates that the neutrons from the nuclear reactor and laboratory environment with energy less than 10 MeV [59] that also shows in figure 6, can not contribute to the backgrounds.
After processing the MC data, and passing the LXe-veto, single-scatter and FV selections, the CRNs-induced NR background rate is estimated to be (6.0 0.6) in keVnr ROI of CENS searching, with the recoil energy spectrum shows in figure 7 (top). In addition, the secondaries of CRNs will also induce ER background in LXe, with a rate of (3.9 0.3), which is shown in figure 7 (bottom).
@@ V.2 Background induced by cosmic muons
The RELICS detector will operate without overburden. Hence, special attention has to be paid to the muon-related background. Even though muons can be tagged with 99% efficiency by the PS veto detector, the neutron background produced by muon interacting with the detector materials is still one of the main background components. Dedicated simulations are performed to evaluate the muon-induced backgrounds. The energy and angular distributions of muons follow the Shukla model [60]. The energy deposition in LXe that is caused by muon-induced fast neutron and related gammas are recorded by RelicsSim, and then filtered with LXe-veto, single-scatter and FV selections. Assuming a 99% muon tagging efficiency in 300 veto time, the muon-induced NRs and ERs are (0.06 0.01) in keVnr and (28.0 2) in keVer, respectively. The corresponding spectra are shown in figure 7.
@@ V.3 Neutrons and gammas from the detector materials
S2 width selection: As shown in figure 8, a large amount of background events are located close to the liquid-gas interface. These backgrounds are mainly from the radioactivity of PMTs and can not be suppressed with the selection in “z” in an “S2-only” analysis. However and as demonstrated in [37], the pulse shape of S2 signals can discriminate background happening at the liquid-gas interface to signals distributed uniformly in depth “z”. A toy MC was conducted to consider the drift and diffusion properties of the electron clouds in LXe to optimize such a selection. We simulate S2s and their pulse shape parameters with “z” distributions shown in figure 8 with energy in the CENS ROI. The S2s are simulated via RelicsAPT where the NEST model is incorporated. The S2 pulse shape is simulated using the electron diffusion model , where is the longitudinal diffusion coefficient, is the electron drift velocity, is the width of single electron waveform [62]. In our simulation. = 16 cm2/s, = 0.174 cm/s, and = 0.19s.
@@ V.4.2 Cosmogenic isotopes and
The RELICS experiment, which is to be operated without overburden, is inevitably exposed to cosmic radiation. We consider cosmogenic radioactive isotopes through the activation and spallation of xenon atoms. In particular, and are known cosmogenic backgrounds in LXeTPCs [65, 34] due to their characteristic x-rays or Auger electrons through electron capture process and their relatively long half-life of 36.4 and 35.0 days, respectively.
Natural xenon does not contain . can be produced from when cosmogenic muons kick off a neutron in the nucleus with a cross-section of (2.74 0.4) cm2 [66]. Additionally, we consider another production channel through the neutron capture on with an upper limit cross-section of 0.7cm2 at 95% C.L [66]. The production rate of in LXe is calculated based on its production cross-section, the muon flux [60] and thermal neutron flux [67]. The decays to via electron capture (EC) process, with a half-life of 36.4 days. We assume that the production and decay of reach equilibrium before deploying the RELICS detector inside the shielding. The estimated equilibrium decay rate is less than 9.5 atoms/day/kg. The shielding system of RELICS can attenuate neutrons efficiently, and the equilibrium decay rate can decrease to 0.79 atoms/day/kg after 300 days.
@@
We consider three possible mechanisms for production in xenon, following the approaches in [68]. Firstly, the extremely low content of in the atmosphere, which is generated by spallation of through (n, 4n), neutron capture of , or cosmic bombardment of calcium-containing soils, via (n, ). This mechanism only needs to be considered in the context of air leaks of an LXe detector, which is strictly limited to below 0.1 liter per year to ensure the purity level of LXe. Secondly, the spallation of detector materials, especially iron in steel. However, its contribution is greatly suppressed due to the slow diffusion rate of argon in steel [68]. Lastly, the direct spallation of xenon atoms by high-energy cosmogenic protons and neutrons. Only the last possibility is considered in this work. The semi-empirical formula by Silberberg and Tsao was used to calculate the spallation cross-section [69]. The CRY is used to calculate the proton spectrum. Gordon’s neutron spectrum [70] is adopted for cosmogenic neutron estimation. The production rate of this process is estimated to be 0.024 atoms/kg/day with an uncertainty of 100% [68]. The produced decays to the ground state of via EC and subsequently deposits energy via x-rays or Auger electrons. The total energy (relative intensity) for K, L, and M shells are 2.82 keV (90.2%), 0.270 keV (8.9%), and 0.018 keV (0.9%), respectively. As a results, the background rate of in the CENS ROI is below kg-1day-1 and completely negligible for CENS search.